# Adaptive Voxel-Weighted Loss Using L1 Norms in Deep Neural Networks for Detection and Segmentation of Prostate Cancer Lesions in PET/CT Images

### v1

@@ Abstract
This study proposes a new loss function for deep neural networks, L1-weighted Dice Focal Loss (L1DFL), that leverages L1 norms for adaptive weighting of voxels based on their classification difficulty, towards automated detection and segmentation of metastatic prostate cancer lesions in PET/CT scans. We obtained 380 prostate-specific membrane antigen (PSMA) [18F]DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence metastatic prostate cancer. We trained two deep neural networks with 3D convolutional neural network backbone — Attention U-Net and SegResNet — and concatenated the PET and CT volumes channel-wise as input. The performance of our custom loss function was evaluated against the Dice and Dice Focal Loss functions. For clinical significance, we considered a detected region of interest (ROI) as a true positive if at least the voxel with the maximum standardized uptake value falls within the ROI. We assessed the models’ performance based on the number of lesions in an image, tumor volume, activity and extent of lesion spread. The L1DFL outperformed the comparative loss functions by at least $13\%$ on the test set. In addition, the F1 scores of the Dice Loss and the Dice Focal Loss were lower than that of L1DFL by at least $6\%$ and $34\%$, respectively. The Dice Focal Loss yielded more false positives, whereas the Dice Loss was more sensitive to smaller volumes and struggled to segment larger lesions accurately. They also exhibited network-specific variations and yielded declines in segmentation accuracy with increased tumor spread. Our results demonstrate the potential of L1DFL to yield robust segmentation of metastatic prostate cancer lesions in PSMA PET/CT images. The results further highlight potential complexities arising from the variations in lesion characteristics that may influence automated prostate cancer tumor detection and segmentation. The code is publicly available at: https://github.com/ObedDzik/pca_segment.git.
@@
[a]organization=School of Biomedical Engineering,addressline=University of British Columbia, city=Vancouver, country=Canada
@@
[b]organization=Department of Integrative Oncology,addressline=BC Cancer Research Institute, city=Vancouver, country=Canada
@@
[d]organization=Department of Physics & Astronomy,addressline=University of British Columbia, city=Vancouver, country=Canada
[…]
[e]organization=Department of Radiology,addressline=University of British Columbia, city=Vancouver, country=Canada
[…]
[f]organization=Department of Electrical and Computer Engineering,addressline=University of British Columbia, city=Vancouver, country=Canada
1 Introduction
Semantic segmentation involves mapping every pixel of an image to a class, whether an object or a background [1]. It requires both a local and contextual understanding of the image [2], making it particularly difficult since neighboring pixels are highly related; thus, information on the spatial context and the boundary’s smoothness is essential in a segmentation model [3]. Particularly in medical image analysis, higher accuracy levels are required for tumor segmentation since false detections may significantly impact patient management [4]. However, medical images are fraught with an imbalance in the proportion of voxels belonging to the regions of interest (ROIs) compared to the background making a balance in precision and sensitivity in automated segmentation a challenging task.
Whole body positron emission tomography (PET) scans of cancer patients are an example of such medical images where voxels corresponding to the foreground class are overwhelmingly outnumbered by background voxels. These images are characterized by high variability in lesion size, shape, and intensity, as well as significant heterogeneity in tracer uptake [5]. In the case of metastatic prostate cancer (mPCa) images, lesions can range from small, low-contrast regions barely distinguishable from surrounding tissues to large, high-uptake areas with well-defined boundaries, and may occupy different anatomical sites [6]. In localizing these lesions, the prostate-specific membrane antigen (PSMA) is often targeted due to its high expression by tumor cells. However, PSMA expression is not exclusive to prostate cancer lesions; it is also found in other tissues, including salivary glands, kidneys, and the small intestine, as well as in certain non-prostatic malignancies and inflammatory processes [7]. This broad expression can lead to radiotracer uptake in non-cancerous or non-prostatic regions challenging the specificity of computer-aided detection algorithms. Thus, the challenge in artificial intelligence (AI) based lesion detection and segmentation in PET scans of metastatic prostate cancer lesions is avoiding false-positive segmentation of anatomical structures and other artifacts while maintaining accurate lesion detection.
With most of the automated approaches, improved sensitivity of mPCa lesions often comes at the cost of specificity, yet both extremes pose potential risks. One commonly used strategy for limiting false positive rate is standardized uptake value (SUV) thresholding, which involves defining all voxels with SUV greater than some value, typically considering the background uptake of the liver to be foreground and all other voxels as background [8]. This method may be included as part of the automated segmentation process or as a post-processing step and has been shown to reduce false positive counts [9]. Yet, its major drawback is the potential of having a reduced sensitivity for lesions that exhibit lower avidity for the radiotracer, or in patients with extensive liver metastases [10].
Loss function design is another strategy that has received much attention and has led to an improvement in the performance of AI-based segmentation methods. The widely used loss function is the Dice Loss. It is based on the Dice Similarity Coefficient (DSC) and penalizes pixel-wise discrepancies between the predicted maps and their corresponding ground truth for each class [11]. A model trained using the Dice Loss objective maximizes DSC inherently incorporating a balance in sensitivity and specificity [12]. Nevertheless, the Dice Loss is not well-suited for handling another type of imbalance caused by a difference in the difficulty in classifying pixels [13]. Within a class, be it the minority foreground class or majority background class, some pixels might be easier to classify, with models achieving greater than $50\%$ probability of correctly classifying such pixels [13]. However, others may be more challenging to predict, yielding lower probabilities [13] and are therefore designated hard to classify. Without a focusing strategy implemented, the Dice Loss can be easily overwhelmed by the easy-to-classify samples and may perform poorly on the harder ones [14]. Besides, it exhibits an inconsistent performance when segmenting multiple lesions of varying sizes in an image [15].
These limitations have prompted the development of alternative loss functions. Distribution-based loss functions like Focal and Cross-Entropy losses are frequently used alternatives due to their dynamic focusing ability. They are often combined with the Dice Loss and its variants to form compound loss [11] since the losses based on overlap measurements are more robust for medical segmentation tasks [16]. With an additional ability to control the model’s focus on detecting ROIs, these combo losses generalize well [11, 17]. The challenge, however, remains on the balance of sensitivity and specificity as these compound losses tend to lose the Dice Loss component, potentially leading to more false positive segmentation [18]. Thus, various regularizations and additional constraints are required to improve model accuracy for a medical lesion segmentation task [19].
This work proposes a new loss function, L1-weighted Dice Focal Loss (L1DFL), with a dynamic weighting strategy based on L1 norms of predicted probabilities and ground truth labels. Using L1 norms, the loss function considers the difficulty of each sample and the prevalence of similar samples in the dataset. In medical image segmentation tasks, datasets are often imbalanced (disproportionately higher number of background voxels than ROIs) and dominated by easy samples (voxels that are easily classified correctly); the L1 norms of these easy samples tend to approach zero. Our technique focuses on down-weighting both easy samples and those with high L1 norm counts, which are likely to represent the background. We demonstrate the efficiency of L1DFL by experiments on [18F]DCFPyL PET/CT scans of patients with biochemical recurrence metastatic prostate cancer, for which lesions have varied sizes, primarily small, with low uptake values.
The main contributions of this work are as follows: (i) we introduce a novel loss function, L1-weighted Dice Focal Loss (L1DFL), incorporating a dynamic weighting strategy; (ii) we evaluate the performance of L1DFL, Dice Loss, and Dice Focal Loss functions in handling false positives and false negative rates; (iii) we compare the segmentation performance of the loss functions based on different lesion scenarios: single-lesion, and multiple lesions. These scenarios are based on the number of lesions in an image; (iv) we evaluate the performance of these loss functions on different clinical metrics like molecular tumor volumes and the extent of lesion spread in the body; and (v) we assess performance using two network architectures, Attention U-Net [20] and SegResNet [21].
2 Materials and Methods
2.1 Dataset
We analyzed 380 prostate-specific membrane antigen (PSMA) [18F]DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence metastatic prostate cancer as part of an ongoing clinical trial (NCT02899312) with informed written consent obtained from the subjects. The image data collection was approved by the University of British Columbia – BC Cancer Ethics Board. The inclusion criteria for the patients were: (1) histologically proven prostate cancer with biochemical recurrence after initial curative therapy with radical prostatectomy, with a PSA $>$ 0.4 ng/mL and an additional PSA measurement confirming an increase; and (2) histologically proven prostate cancer with biochemical recurrence after initial curative radiotherapy, with a PSA $>$ 2 ng/mL after therapy, with five or fewer lesions identified on [18F]DCFPyL PET/CT [22].
Each patient received an average dose of 350 MBq [18F]DCFPyL after fasting for four hours, with dose adjustments based on body weight. Imaging was conducted 120 minutes after injection using a GE Discovery 600 or 690 scanner (GE Healthcare, USA). A non-contrast-enhanced CT scan was acquired for localization and attenuation correction (120 kV, automatic mA adjustment from 30–200 mA, and a noise index of 20). Following CT acquisition, a whole-body PET scan was conducted at 2–4 minutes per bed position, depending on the participant’s size, with reconstruction performed using an ordered subset expectation maximization algorithm and point-spread function modeling. Each transaxial PET image had a matrix resolution of 192 × 192 pixels, with a physical pixel size of $3.64\text{mm}^{2}$. An expert nuclear medicine physician performed manual segmentation on all active lesions [22].
There were a total of 684 prostate cancer (PCa) lesions across the entire dataset with a mean active lesion volume of 6.68 $\pm$ 10.20 ml. The lesions varied in uptake with the average maximum standardized uptake value (SUVmax) and mean standardized uptake value (SUVmean) being 12.65$\pm$14.46 and 4.62 $\pm$ 3.88, respectively.
2.2 Ground Truth Annotation
The ground truth segmentations were performed by a board-certified nuclear medicine physician from the BC Cancer Research Institute. All lesions were annotated using the PET Edge tool, a semi-automated gradient-based segmentation tool, and contours refined using the 3D Brush tool. Both tools are available in the MIM workstation (MIM software, Ohio, USA). These segmentations served as the ground truth labels for training the models.
@@ 2.3 Data Preprocessing and Augmentation
2.4 Architectures
In this work, we implemented and studied two convolutional neural network architectures: Attention U-Net [20] and SegResNet [21]. These networks, built on top of a 3D-CNN backbone, comprise encoder and decoder blocks. The SegResNet uses residual blocks with skip connections for feature propagation and maintaining gradient stability during training [21]. The encoder path contained blocks with the following number of layers: 1, 2, 2, and 4, allowing for increasing levels of feature extraction at different resolutions. The initial filter size was set to 16 channels, then doubled through each downsampling step to capture progressively higher-level features. Each convolutional layer employs a ReLU activation function, and Group Normalization [23]. The decoder path, containing one layer in each block, gradually reconstructs the segmented output while leveraging skip connections to preserve spatial details from the encoder. The input and output channels were set to 2, accommodating multi-channel input images and dual-class segmented outputs.
On the other hand, Attention U-Net extends the U-Net architecture to include attention gates along the encoder-decoder path, allowing the network to selectively pay attention to essential features [20]. The encoder path consisted of five layers, with each layer increasing in channel depth from 16, 32, 64, 128, to 256 channels, and each layer downsampled by a stride of 2. These attention gates guide the network in emphasizing regions that contribute significantly to the segmentation task. Each convolutional layer employs a ReLU activation function, and Batch Normalization as in the original implementation [20]. The output layer matched the input with two channels for multi-modal inputs and dual-class output, representing background and foreground classes.
2.5 Model Training
We divided the dataset into training, validation, and testing sets containing 258, 65, and 57 samples, respectively. The training objective was to minimize the loss functions in the training set. We used AdamW optimizer with weight decay of $10^{-5}$ to optimize the loss functions. We adopted a cosine annealing scheduler to decay the learning rate from $2\times 10^{-4}$ to zero for 1000 epochs. The loss was first computed for each batch within an epoch and the overall loss for an epoch was calculated by taking a mean over all the batch losses. The model with the highest mean DSC in the validation phase was chosen for further evaluation of the test set. The code for the implementations is made available at: https://github.com/ObedDzik/pca_segment.git
@@ 2.6 Loss Functions
Dice Loss (DL):
The Dice Loss maximizes the overlap between the predicted segmentation and the ground truth. In this work, for a specific cubic patch of an image, the Dice Loss for a batch was computed as,
@@
$\mathcal{L}_{\text{Focal}}=-\frac{1}{2}\sum_{c\in\{0,1\}}\sum_{i}\alpha_{c}(1-p_{i}(c))^{\gamma}\log(p_{i}(c))$ (2)
where $p_{i}$ is the predicted probability for a given class $c$ of a voxel $i$, $\alpha$ is a factor defining the balance between background and foreground classes in a binary segmentation task, and $\gamma$ is the focusing parameter that adjusts the rate at which easy examples are down-weighted.
Dice Focal Loss (DFL):
The Dice Focal Loss [25] combines the Dice Loss and the Focal Loss forming a compound loss function:
@@
For the Dice Focal Loss formulation, we set $\gamma=2$ and $\alpha=1$ in all our experiments similar to the implementation reported in [26].
L1-weighted Dice Focal Loss (L1DFL):
The L1-weighted Dice Focal Loss (L1DFL) applies a weighting strategy to the Dice Loss (DL) based on the L1 norm between the predicted probabilities and the ground truth labels and combines it with the Focal Loss (FL). The Dice Loss generally has two variants, one with squared terms in the denominator [27] and one without [12]. In L1DFL, we employ the Dice Loss with the squared denominator which was preferred over the non-squared version following its performance in [28]. We describe the mathematical formulation of the weighting strategy next.
@@
where $n=\lceil{\frac{1}{\Gamma}}+1\rceil$ is the total number of bins. For instance, with a bin width of $\Gamma=0.1$ and a total of $n$ = 11 bins, $B_{k}$ takes values at regular intervals in the range [0,1], such that for $k=0,1,2,\dots,10$, bin centers ($B_{k}$) = $[0,0.1,0.2,\dots,1.0]$.
@@
The density $\mathcal{D}(B_{k})$ represents the concentration of samples around the bin center $B_{k}$. For $B_{k}$ values close to 0, the L1 norm values ($\Delta_{i}$) are also close to 00, indicating easy samples where the predictions align closely with the ground truth. In contrast, bins with centers near 1 represent L1 norms closer to 1, indicating hard-to-classify samples where predictions deviate significantly from the ground truth. Thus, based on the norm density, we can evaluate how common or rare a particular prediction difficulty is, allowing the weighting to be adjusted accordingly. This adjustment ensures that less frequent and difficult samples, which often correspond to foreground samples in an imbalanced scenario receive higher weights than the easy-to-classify samples. The weight for each bin is then calculated as:
@@
These weights are applied to the numerator and the denominator of the Dice Loss to account for the imbalances caused by the variations in $\Delta$
@@
where $w_{i}=w(B_{k})$ if $\Delta_{i}$ belongs to bin $B_{k}$. If the L1 norms of the examples are uniformly distributed, the density $\mathcal{D}(B_{k})$ will be the same for all bins, resulting in equal weights $w(B_{k})$ for all bins. In this case, the weighted Dice Loss, ($L_{\text{wDice}}$), will reduce to the standard Dice Loss.
@@
[…]
We illustrate the dynamic weighting strategy of L1DFL below. For a given 4x4 matrix, Figure 1 highlights how L1DFL downweights contributions from the background class and focuses on regions of classification difficulty. Specifically, the figure shows that higher weights are applied for voxels with L1 norm ($\Delta$) close to 1, but for voxels with L1 norms close to 0, representing that those voxels are easily classified, lower weights are applied. Thus, regions corresponding to false positives and negatives receive much attention from the model. In our experiments, we empirically selected a constant bin width, $\Gamma$, of 0.1 and $\gamma$ of 2 for the focal loss component of the loss function.
Figure 1: An illustration of the dynamic weighting strategy in L1 weighted Dice Focal Loss (L1DFL). (a) shows the ground truth (G), where 1 indicates lesion areas and 0 indicates background. (b) displays the predicted probability map, where green regions denote false positive predictions and red regions represent ROIs. (c) shows the L1 norm values between the predicted probabilities and ground truth, highlighting the error magnitude for each voxel. Higher values imply higher classification difficulty. (d) presents the density, $\mathcal{D}$, which reflects the frequency distribution of each L1 norm value, capturing the difficulty of correctly classifying each voxel. Lower values imply higher classification difficulty. (e) depicts the final weights assigned to each voxel, where higher weights focus on rare, misclassified regions while down weighting well-classified background areas. This weighting strategy reduces the influence of abundant, correctly predicted voxels and increases the focus on difficult, misclassified lesions, thereby enhancing the model’s focus on challenging samples.

### v2

@@ Abstract
Background and Objective: Accurate automated detection of recurrent prostate cancer in PSMA PET/CT scans is challenging due to heterogeneous lesion size, activity, anatomical location, and intra- and inter-class imbalances. Conventional deep learning loss functions often produce suboptimal optimization, as gradients are dominated by easy background voxels or extreme outliers. To address this, we propose L1-weighted Dice Focal Loss (L1DFL), which harmonizes gradient magnitudes across voxels using L1 norms to adaptively weight samples based on classification difficulty, resulting in well-calibrated predictions with a bimodal separation between correct and incorrect predictions.
Methods: We trained three 3D convolutional networks (Attention U-Net, SegResNet, U-Net) and a transformer-based UNETR model on 380 PSMA PET/CT scans. PET and CT volumes were concatenated as input. A finetuned medical foundation model (SAM-Med3D) was also evaluated. L1DFL was compared against Dice Loss (DL) and Dice Focal Loss (DFL) using metrics including Dice Similarity Coefficient (DSC), F1 score, and lesion-wise detection performance, with statistical significance assessed via Wilcoxon signed-rank tests with Bonferroni correction ($\alpha=0.003$).
Results: Across architectures, L1DFL consistently outperformed DL and DFL, achieving at least a $4\%$ improvement in DSC. F1 scores were higher by $\geq 6\%$ and $\geq 26\%$ compared to DL and DFL, respectively. While DFL produced more false positives and DL struggled with larger lesions, L1DFL achieved balanced detection, minimizing false detections while maintaining high true positive rates. The gradient harmonization mechanism ensured robustness across varying lesion sizes, volumes, and spread.
Conclusion: L1DFL improves segmentation of recurrent prostate cancer lesions in PSMA PET/CT by harmonizing gradients. It produces well-calibrated predictions with clear separation of correct and incorrect cases, and balanced detection across lesion sizes while maintaining consistent performance across multiple architectures.
@@
[a]organization=University of British Columbia,city=Vancouver, country=Canada
@@
[b]organization=BC Cancer Research Institute,city=Vancouver, country=Canada
@@
{highlights}
Besides uptake, tumor characteristics such as volume and dissemination limit automated detection performance.
The proposed loss function demonstrates strong calibration, producing a clear separation between correct and incorrect predictions.
It dynamically distributes gradient magnitudes across samples, avoiding overemphasis on outliers or overly easy examples.
It outperforms the Dice and Dice Focal Losses by at least $4\%$ based on Dice Score.
It outperforms Dice Loss and Dice Focal Losses by $6\%$ and $26\%$, respectively, on F1 score.
1 Introduction
Semantic segmentation assigns a label to every pixel or voxel in an image [9], requiring both local feature extraction and broader contextual understanding [5], since spatial coherence and boundary smoothness are critical [34]. In medical imaging, segmentation accuracy is particularly critical, as false detections or missed lesions can significantly affect diagnosis, staging, and treatment planning [28]. These difficulties are compounded by severe class imbalance, where pathological voxels constitute only a small fraction of the image volume.
In whole-body oncological PET, lesion voxels are heavily outnumbered by background, and lesions vary widely in size, shape, contrast, and tracer uptake [15]. Metastatic prostate cancer (mPCa) amplifies these challenges further, with lesions distributed across multiple anatomical sites [10]. PSMA PET imaging adds another layer of complexity, as PSMA expression appears in normal organs and inflammatory processes as well as malignant tissue [19]. AI-based detection systems must therefore balance sensitivity to small or low-uptake lesions against specificity to avoid false positives from physiological uptake.
Most automated approaches struggle with this trade-off. Techniques such as standardized uptake value (SUV) thresholding are commonly employed post hoc to suppress false positives by classifying voxels above a predefined threshold as foreground [11]. While effective in reducing spurious detections, such strategies often sacrifice sensitivity for lesions with low tracer avidity or lesions in patients with atypical uptake patterns [23]. More sophisticated methods have attacked the problem at the input or representation level: semi-automatic pipelines with bone masking and SUV thresholding [12], nnU-Net-based whole-body frameworks [16, 17], anatomical organ priors [23], 2D diffusion models on MIP projections [37], and hybrid MIP-based frameworks [7]. Feature-clipping normalization [3] and 2.5D pipelines [44] have also been explored. While each has yielded incremental gains, false positives and missed small or ambiguous lesions remain persistent failure modes, and critically, none directly addresses how voxel-level errors are weighted during optimization.
In PET lesion segmentation, not all voxels contribute equally to learning. The majority of background voxels are trivially classified, while a small subset of foreground and boundary voxels is consistently misclassified or uncertain. Standard losses do not distinguish between these difficulty levels. Region-based losses such as Dice Loss (DL) handle inter-class imbalance through overlap-based formulation [8, 42] but treat all voxels within a class equally [43]. Distribution-based losses like Focal Loss [24] and TopK loss [39] address difficulty-based weighting but rely on fixed hyperparameters that limit adaptability to mini-batch distributions [26], and improper weighting can compromise sensitivity-specificity balance [41]. Compound losses such as Dice Focal Loss [45] and Combo Loss [36] leverage complementary strengths but inherit the same static weighting limitations.
A more principled solution is gradient harmonization [22], which adaptively weights samples based on gradient density rather than fixed hyperparameters, preventing easy examples from overwhelming the contribution of hard minority cases. Liu et al. [26] extended this to medical image segmentation via knee MRI, demonstrating its value for handling intra-class imbalance. However, this approach has not yet been explored for PET lesion segmentation, where the gradient density problem is especially acute.
We therefore propose L1-weighted Dice Focal Loss (L1DFL), which integrates gradient harmonization into the Dice Focal Loss framework. Unlike prior approaches that bin samples by computed gradient magnitudes [22, 26], L1DFL uses the L1 norm between predicted probabilities and ground truth labels as a canonical proxy for voxel-level classification difficulty [21, 33, 25]. Voxels in high-density difficulty regions are down-weighted while those in low-density regions are up-weighted, distributing gradient contributions uniformly across the difficulty spectrum and avoiding domination by any single difficulty level during optimization.
The main contributions of this work are: (i) a novel loss function, L1DFL, incorporating gradient harmonization; (ii) evaluation of L1DFL against DL and DFL on false positive and false negative rates; (iii) comparison across single-lesion and multiple-lesion scenarios; (iv) assessment on clinical metrics including molecular tumor volume and lesion spread; and (v) benchmarking across five models: Attention U-Net [32], SegResNet [30], U-Net [6], UNETR [14], and SAM-Med3D [38].
2 Materials and Methods
2.1 Dataset and Ground Truth Annotation
We analyzed 380 [18F]DCFPyL PET/CT scans of patients diagnosed with biochemical recurrence of prostate cancer as part of an ongoing clinical trial (NCT02899312) with informed written consent obtained from the participants. The image data collection was approved by the University of British Columbia – BC Cancer Ethics Board. The inclusion criteria for the patients were: (1) histologically proven prostate cancer with biochemical recurrence after initial curative therapy with radical prostatectomy, with a PSA >> 0.4 ng/mL and an additional PSA measurement confirming an increase; and (2) histologically proven prostate cancer with biochemical recurrence after initial curative radiotherapy, with a PSA >> 2 ng/mL after therapy, with five or fewer lesions identified on [18F]DCFPyL PET/CT [13].
Patients received an average of 350 MBq of [18F]DCFPyL following a four-hour fast, with dose adjustments based on body weight. PET/CT imaging was performed 120 minutes post-injection using GE Discovery 600 or 690 scanners (GE Healthcare, USA). A non-contrast CT scan was acquired for attenuation correction and anatomical localization (120 kV, automated mA modulation), followed by a whole-body PET acquisition lasting 2-4 minutes per bed position. PET images were reconstructed using ordered subset expectation maximization with point-spread function modeling, producing transaxial images with a matrix resolution of $192\times 192$ and an in-plane pixel spacing of $3.64~\text{mm}$ x $3.64~\text{mm}^{2}$ and a slice thickness of $3.27~\text{mm}$ [13].
@@ 2.2 Data Preprocessing and Augmentation
2.3 Implementation Details
Architectures:
We evaluated five volumetric segmentation architectures: SegResNet [30], Attention U-Net [32], 3D U-Net [6], UNETR [14], and SAM-Med3D [38]. SegResNet, Attention U-Net, and 3D U-Net were implemented as convolutional encoder–decoder networks with multi-channel inputs and two-class outputs. SegResNet employed encoder depths of {1, 2, 2, 4} with an initial channel size of 16, doubling at each downsampling stage, and used Group Normalization and ReLU activations. Attention U-Net consisted of five encoder levels with channel sizes {16, 32, 64, 128, 256}, Batch Normalization, and attention gates as originally proposed. The 3D U-Net comprised six levels with channels increasing from 16 to 512, residual blocks with Batch Normalization, and a symmetric decoder with transposed convolutions.
UNETR and SAM-Med3D represent transformer-based and foundation-model approaches. UNETR used a vision transformer encoder with hidden size 256, MLP dimension 1024, and four attention heads, convolutional patch embeddings, and a U-shaped convolutional decoder. SAM-Med3D is a fully three-dimensional extension of the Segment Anything Model designed for volumetric medical image segmentation. It follows the original SAM design, consisting of a 3D image encoder, a 3D prompt encoder, and a mask decoder, with all 2D operations replaced by 3D counterparts, including 3D convolutions, normalization layers, and positional encodings [38]. We used the publicly released SAM-Med3D pretrained weights and fine-tuned the model using our proposed and comparative loss functions.
Model Training:
We divided the dataset into training, validation, and testing sets containing 258, 65, and 57 samples, respectively. The training objective was to minimize the loss functions in the training set. We used AdamW optimizer with a weight decay of $10^{-5}$ to optimize the loss functions. We adopted a cosine annealing scheduler to decay the learning rate from $2\times 10^{-4}$ to zero for 1000 epochs. The loss was first computed for each batch within an epoch, and the overall loss for an epoch was calculated by taking a mean over all the batch losses. The model with the highest mean DSC in the validation phase was chosen for further evaluation of the test set.
For SAM-Med3D, we evaluated both zero-shot performance and fine-tuning. Fine-tuning was conducted using the original loss function as well as comparative loss functions for evaluation. Training used a batch size of 2, an initial learning rate of $8\times 10^{-6}$ and 150 epochs. Optimization employed AdamW with a linear learning rate scheduler (decay factor 0.1).
@@ 2.4 Loss Functions
Dice Loss:
The DL maximizes the overlap between the predicted segmentation and the ground truth. In this work, for a specific cubic patch of an image, the DL for a batch was computed as,
@@
$\mathcal{L}_{\text{Focal}}=-\sum_{c\in\{0,1\}}\sum_{i}g_{i}(c)\alpha_{c}(1-p_{i}(c))^{\gamma}\log(p_{i}(c))$ (2)
where $\alpha$ is a factor defining the balance between background and foreground classes in a binary segmentation task, and $\gamma$ is the focusing parameter that adjusts the rate at which easy examples are down-weighted.
Dice Focal Loss:
The DFL [45] combines the DL and the Focal Loss forming a compound loss function:
@@
For the DFL formulation, we set the Focal Loss terms $\gamma=2$ and $\alpha=1$ in all our experiments similar to the implementation reported in [31].
L1-weighted Dice Focal Loss:
L1DFL applies a gradient harmonization weighting strategy to the Dice Loss based on both the L1 norm between predicted probabilities and ground truth labels (quantifying voxel-level difficulty) and the density of samples at each difficulty level (defined by bins), then combines this weighted Dice Loss with Focal Loss. The DL generally has two variants, one with squared terms in the denominator [29] and one without [8]. In L1DFL, we employ the DL with the squared denominator which was preferred over the non-squared version following its performance in [26]. We describe the mathematical formulation of the weighting strategy next.
@@
where $n=\lceil{\frac{1}{\Gamma}}+1\rceil$ is the total number of bins. For instance, with a bin width of $\Gamma=0.1$ ($n$ = 11 bins), $B_{k}$ takes values at regular intervals in the range [0,1], such that for $k=0,1,2,\dots,10$, bin centers ($B_{k}$) = $0,0.1,0.2,\dots,1.0]$.
@@
The density $\mathcal{D}(B_{k})$ quantifies how many voxels have L1 norms within the bin centered at $B_{k}$. The bin center represents the difficulty level: $B_{k}\approx 0$ indicates easy samples (predictions closely match ground truth), while $B_{k}\approx 1$ indicates hard samples (predictions deviate significantly from ground truth). Thus, density, $\mathcal{D}(B_{k})$, measures the prevalence of each difficulty level, and high density means many voxels share similar L1 norm values (commonly occurring difficulty), while low density means few voxels fall in that range. The weight for each bin is thus calculated as:
@@
where $N$ is the total number of samples. Consequently, samples in high-density regions receive lower weights, while samples in low-density regions receive higher weights. This ensures that less frequent and difficult samples, which often correspond to small foreground structures in imbalanced scenarios, receive appropriate emphasis without allowing outliers or ambiguous examples to dominate optimization, unlike focal loss approaches that monotonically increase weights with difficulty. These harmonized weights are applied to both the numerator and denominator of the Dice loss to account for imbalances caused by variations in $\Delta$.
@@
where $w_{i}=w(B_{k})$ if $\Delta_{i}$ belongs to bin $B_{k}$. If the L1 norms of the examples are uniformly distributed, the density $\mathcal{D}(B_{k})$ will be the same for all bins, resulting in equal weights $w(B_{k})$ for all bins. In this case, the weighted DL, ($L_{\text{wDice}}$), will reduce to the standard DL.
@@
$\text{L1DFL}=\mathcal{L}_{\text{wDice}}+\mathcal{L}_{\text{Focal}}$ (12)
For our implementation, we empirically selected a constant bin width, $\Gamma$, of 0.1 and $\gamma=2$ for the Focal Loss component of L1DFL. We assess the robustness of the loss function to $\Gamma$ and $\gamma$ hyperparameters, computed over a five-epoch window centered at the epoch with the highest validation Dice. We report the performance in the results section (3.6). We illustrate the dynamic weighting strategy of L1DFL in Figure 1 using a 4×4 matrix.
Figure 1: An illustration of the dynamic weighting strategy in L1DFL.(a) Ground truth (G) with lesion areas (1) and background (0). (b) Predicted probability map showing false positives (green) and regions of interest (red). (c) L1 norm values computed as $|\hat{y}-y|$ for each voxel, representing prediction error magnitude. Higher L1 norm values indicate greater classification difficulty. (d) Density $\mathcal{D}$, showing the frequency of each voxel’s difficulty level across all voxels. A difficulty level is defined by a range of L1 norm values (bin), and density measures how many voxels fall within that bin. (e) Final weights ($w=N/\mathcal{D}$) assigned through gradient harmonization. Voxels at high-density difficulty levels receive lower weights, while voxels at low-density difficulty levels receive higher weights. This prevents gradient concentration on any single difficulty level, distributing optimization effort more uniformly across the difficulty spectrum.
2.5 Empirical Loss Analysis
To empirically validate the gradient harmonization mechanism of L1DFL and compare it against baseline loss functions, we conducted two complementary analyses on converged models.
Prediction Distribution Analysis:
We evaluated the calibration properties and uncertainty quantification capabilities of each loss function by analyzing the distribution of prediction entropy and confidence scores. For each voxel in the validation set, we computed the prediction entropy as $H=-\sum_{c}p_{c}\log p_{c}$ where $p_{c}$ is the predicted probability for class $c$, and the prediction confidence as $\max_{c}p_{c}$. We stratified these distributions by prediction correctness (correct vs. incorrect predictions) to assess whether each loss function produces well-calibrated uncertainty estimates. The calibration gap, defined as the difference in mean confidence between correct and incorrect predictions, quantifies how well the model’s confidence aligns with its actual accuracy, a larger gap indicates better calibration.
Sample Weighting Analysis:
To directly demonstrate the gradient harmonization effect, we analyzed how each loss function weights samples of varying difficulty. For each test sample, we computed: (1) sample difficulty as the mean absolute difference between predictions and targets $d=|\hat{y}-y|$, representing the inherent difficulty of the prediction, and (2) effective sample weight as the $L_{2}$ norm of the loss gradient with respect to the model outputs $w=\|\nabla_{\hat{y}}\mathcal{L}\|_{2}$, representing how much each sample contributes to the gradient during optimization. We computed the Pearson correlation coefficient $\rho(d,w)$ between difficulty and weight across all samples. A lower correlation indicates that the loss function assigns more balanced weights across the difficulty spectrum, preventing gradient concentration on outliers. We also visualized the relationship between difficulty and weight by binning samples into 50 difficulty levels and computing mean weights per sample for each bin. This analysis reveals whether a loss function over-emphasizes hard samples (high $\rho$) or achieves balanced weighting preventing outlier dominance.
Both analyses were performed on a converged SegResNet model trained with the different loss functions using the validation set, with sliding window inference for computational efficiency.